Hello, i need help on this question please i have tried to solve but just can't. Thank you
y/2<-3

Answer by lenny460(1073)

Hello, I need help on this question please.
I have tried to solve but just can't.
Thank you
y/2<-3

y/2 < -3
Multiply each side of the inequality by 2
y/2 * 2 < -3 * 2
y/2 * 2 = y
-3 * 2 = -6
Answer:
y < -6


Lennox Obuong
Algebra Tutor